Calculate the missing statistics in each line. INCLUDE YOUR UNITS FOR EACH VALUE

Subject           Heart Rate                  Stroke volume                    CO          

1                      50                                            100                              ____(A)______

2                      30                                            80                                ____(B)______

3                      60                                __    __(C)______                               7200

4                      _(D)_                                       45                                            900
Answer #1

1. Cardiac output = Heart rate \times stroke volume

                             = 50 \times 100

(A) is 5000

2. Cardiac output = Heart rate \times stroke volume

                             = 30 \times 80

(B) is 2400

3. Stroke volume = cardiac output \div heart rate

                            = 7200 \div 60

(C) is 120

4. Heart rate = cardiac output \div stroke volume

                     = 900 \div 45

(D) is 20